# Robust Preference for Dynamical Dark Energy in DESI BAO and SN Measurements

### v1

@@ IV.2 Results for the Exponential parametrization
Figure 2: Exponential parameterization (6) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
@@
$w_{0}$ $-0.50\pm 0.27$ $-0.820\pm 0.064$ $-0.804^{+0.045}_{-0.051}$
$w_{a}$ $-1.40^{+0.75}_{-0.62}$ $-0.77\pm 0.28$ $-0.71^{+0.23}_{-0.19}$
$\Omega_{\mathrm{m}}$ $0.352\pm 0.035$ $0.3088\pm 0.0070$ $0.3157\pm 0.0066$
$\sigma_{8}$ $0.788^{+0.026}_{-0.029}$ $0.8186\pm 0.0094$ $0.8150\pm 0.0092$
$S_{8}$ $0.852\pm 0.017$ $0.8304\pm 0.0097$ $0.8360\pm 0.0098$
@@
Table 3: Exponential Parametrization (6) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{exp}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.3 Results for the JBP parametrization
@@
Finally, we replace PantheonPlus with DESY5 SN measurements. In this case, we obtain $w_{0}=-0.641^{+0.095}_{-0.067}$ and $w_{a}=-2.12^{+0.38}_{-0.68}$, confirming that the evidence of DDE becomes much more pronounced with DESY5. This evidence reaches a statistical significance $\gtrsim 3\sigma$. Having said that, comparing Fig. 3 with the respective triangular plots of the other parameterizations, we notice that for this model, the uncertainties remain much broader, especially for the parameters describing DE EoS. This suggests that this parameterization might not be the ideal fitting model for current observations, hinting at some possible internal tension compared to the other cases. This concern is confirmed when comparing the difference between the best-fit $\chi^{2}$ obtained within each DDE model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM. Indeed, this model consistently leads to the smallest improvement over $\Lambda$CDM across all data and DDE models.
Figure 3: JBP parametrization (7) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
@@
Table 4: JBP parametrization (7) – 68% CL constraints and 95% CL upper limits on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{JBP}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.4 Results for the Logarithmic parametrization
@@
Figure 4: Logarithmic parametrization (8) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
@@
Table 5: Logarithmic parametrization (8) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{log}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.5 Results for the BA parametrization
@@
Interestingly, when comparing the difference between the best-fit $\chi^{2}$ obtained within each DDE model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM, this model consistently leads to the most significant improvement over $\Lambda$CDM across all three data combinations analyzed. It performs better than the CPL parameterization, as seen by comparing the last line in Tab.2 and the last line in Tab.6.
Figure 5: BA parametrization (9) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
@@
Table 6: BA parametrization (9) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{BA}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
V Discussions and Conclusions
@@
Once more, all models exhibit similar trends: for Planck+DESI, we consistently observe an improvement in the fit over $\Lambda$CDM, with $\Delta\chi^{2}$ ranging from $-5.6$ to $-8.7$, depending on the specific model. This improvement in the fit is further enhanced when PantheonPlus SN measurements are included ($\Delta\chi^{2}$ ranges from $-6.4$ to $-9.4$) and is substantially increased – by up to a factor of $\sim 2$ – when adopting DESY5 SN data (in this case, $\Delta\chi^{2}$ ranges from $-14.8$ to $-16.2$). This trend follows the overall preference for dynamical dark energy discussed thus far. Interestingly, the linear CPL parameterization is never the best-fitting model. The equation of state proposed by Barboza-Alcaniz, given by Eq. (9), consistently leads to the most significant improvement in $\Delta\chi^{2}$ over $\Lambda$CDM across all three data combinations analyzed. Conversely, the Jassal-Bagla-Padmanabhan parameterization, given by Eq. (7), shows the smallest improvement in fit compared to $\Lambda$CDM among the models considered. These results suggest that current data are approaching a precision that could enhance our understanding of the physical nature behind the apparent dynamical behavior of dark energy, should future surveys and data releases confirm these findings.
Acknowledgements.
@@
W.G. acknowledges support from the Lancaster–Sheffield Consortium for Fundamental Physics through the Science and Technology Facilities Council (STFC) grant ST/X000621/1. S.P. acknowledges the financial support from the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). E.D.V acknowledges support from the Royal Society through a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from the COST Action CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics” (CosmoVerse), supported by COST (European Cooperation in Science and Technology).

### v2

@@ 4.2 Results for the Exponential parametrization
$w_{0}$ $-0.50\pm 0.27$ $-0.876\pm 0.045$ $-0.804^{+0.045}_{-0.051}$
$w_{a}$ $-1.40^{+0.75}_{-0.62}$ $-0.51^{+0.20}_{-0.17}$ $-0.71^{+0.23}_{-0.19}$
$\Omega_{\mathrm{m}}$ $0.352\pm 0.035$ $0.3088\pm 0.0067$ $0.3157\pm 0.0066$
$\sigma_{8}$ $0.788^{+0.026}_{-0.029}$ $0.8192\pm 0.0098$ $0.8150\pm 0.0092$
$S_{8}$ $0.852\pm 0.017$ $0.8310\pm 0.0095$ $0.8360\pm 0.0098$
@@
Table 3: Exponential Parametrization (2.6) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. 3. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{exp}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
Figure 2: Exponential parameterization (2.6) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
4.3 Results for the JBP parametrization
@@
Table 4: JBP parametrization (2.7) – 68% CL constraints and 95% CL upper limits on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. 3. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{JBP}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
Finally, we replace PantheonPlus with DESY5 SN measurements. In this case, we obtain $w_{0}=-0.641^{+0.095}_{-0.067}$ and $w_{a}=-2.12^{+0.38}_{-0.68}$, confirming that the evidence of DDE becomes much more pronounced with DESY5. This evidence reaches a statistical significance $\gtrsim 3\sigma$. Having said that, comparing Fig. 3 with the respective triangular plots of the other parameterizations, we notice that for this model, the uncertainties remain much broader, especially for the parameters describing DE EoS. This can be explained in terms of the peculiar evolution of the DE EoS obtained in this model. As discussed in detail in Appendix A, among the five models analyzed, the JBP parameterization presents a more articulated phenomenology regarding the evolution of the DE EoS. Due to its quadratic nature in the scale factor, the evolution of the EoS within the JBP parameterization crosses $w=-1$ twice. At low redshift, it behaves similarly to the other parameterizations; however, after the first quintessence-to-phantom transition, $w(z)$ approaches a minimum value around $z\sim 1$ before rising again, leading to a second phantom-to-quintessence crossing at $z\sim 4$. This behavior contrasts with other models, where the EoS remains within the phantom regime. As detailed in Appendix A, the interplay between low and high redshift behaviors results in two different pivot redshifts at low and high $z$. This interplay may contribute to tilting the 2-D probability contours in the $w_{0}$ and $w_{a}$ plane, as shown in Fig. 3 (see also Fig. 6 for comparisons with other models). Additionally, the increased uncertainties at low redshift might suggest that this phenomenology is not ideal for consistently fitting all the data across low and high redshift. This concern is confirmed when comparing the differences between the best-fit $\chi^{2}$ obtained within each DDE model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM. Indeed, this model consistently leads to the smallest improvement over $\Lambda$CDM across all datasets and DDE models. For more details, we refer to Appendix A.
Figure 3: JBP parametrization (2.7) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
4.4 Results for the Logarithmic parametrization
@@
Table 5: Logarithmic parametrization (2.8) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. 3. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{log}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
Figure 4: Logarithmic parametrization (2.8) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
4.5 Results for the BA parametrization
@@
Table 6: BA parametrization (2.9) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. 3. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{BA}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
Figure 5: BA parametrization (2.9) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
Interestingly, when comparing the difference between the best-fit $\chi^{2}$ obtained within each DDE model and the best-fit $\chi^{2}$ obtained within $\Lambda$CDM, this model consistently leads to the most significant improvement over $\Lambda$CDM across all three data combinations analyzed. It performs better than the CPL parameterization, as seen by comparing the last line in Tab. 2 and the last line in Tab. 6. As discussed in Appendix A, when comparing the evolution of the EoS inferred in this model with the other cases analyzed so far, we find that at low redshift it behaves similarly to the CPL parameterization. The most notable differences emerge at $z\gtrsim 1$. In all models, the EoS moves deeply into phantom values (except for the JBP model, where it is compelled to rise back towards quintessence-like values). In contrast, within the BA model, $w(z)$ does not trend towards very negative values at $z\gtrsim 1$. While it remains phantom, it stabilizes on a distinctive, nearly flat plateau.
5 Discussions and Conclusions
@@
Once more, all models exhibit similar trends: for Planck+DESI, we consistently observe an improvement in the fit over $\Lambda$CDM, with $\Delta\chi^{2}$ ranging from $-5.6$ to $-8.7$, depending on the specific model. This improvement in the fit is further enhanced when PantheonPlus SN measurements are included ($\Delta\chi^{2}$ ranges from $-6.4$ to $-9.4$) and is substantially increased – by up to a factor of $\sim 2$ – when adopting DESY5 SN data (in this case, $\Delta\chi^{2}$ ranges from $-14.8$ to $-16.2$). This trend follows the overall preference for dynamical dark energy discussed thus far. Interestingly, the linear CPL parameterization is never the best-fitting model. The equation of state proposed by Barboza-Alcaniz, given by Eq. (2.9), consistently leads to the most significant improvement in $\Delta\chi^{2}$ over $\Lambda$CDM across all three data combinations analyzed. Conversely, the Jassal-Bagla-Padmanabhan parameterization, given by Eq. (2.7), shows the smallest improvement in fit compared to $\Lambda$CDM among the models considered. The only exception is for Planck+DESI+DESY5, where $\Delta\chi^{2}=-16$ indicates a better fit to this dataset compared to the CPL, logarithmic, and exponential parameterizations, although it is still smaller compared to the Barboza-Alcaniz model. For further discussion and physical interpretation of the different phenomenological behaviors of the models analyzed so far, we refer to Appendix A. Specifically, Fig. 7 presents constraints on the evolution of the equation of state with respect to redshift, as inferred from various datasets across all models. Tab. 7 provides constraints on other important properties, such as the pivot redshift, the corresponding values (and uncertainties) of the equation of state, and the epoch of the quintessence-to-phantom transition. Overall, these results support the main conclusions regarding the resilience of the DESI and SN preference for evolving dark energy, while suggesting that current data are approaching a precision that could enhance our understanding of its physical nature, should future surveys and data releases confirm these findings.
Acknowledgments
@@
The authors thank the referee for many important comments that improved the manuscript. W.G. acknowledges support from the Lancaster–Sheffield Consortium for Fundamental Physics through the Science and Technology Facilities Council (STFC) grant ST/X000621/1. S.P. acknowledges the financial support from the Department of Science and Technology (DST), Govt. of India under the Scheme “Fund for Improvement of S&T Infrastructure (FIST)” (File No. SR/FST/MS-I/2019/41). E.D.V acknowledges support from the Royal Society through a Royal Society Dorothy Hodgkin Research Fellowship. This article is based upon work from the COST Action CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics” (CosmoVerse), supported by COST (European Cooperation in Science and Technology).
Appendix A Equation of State, pivot redshift, and phantom crossing across the different models
Model Dataset $z_{p}$ $w(z_{p})$ $z_{c}$
CPL Planck+DESI+PantheonPlus $0.27$ $-0.982\pm 0.028$ $0.31^{+0.08}_{-0.06}$
Planck+DESI+DESY5 $0.25$ $-0.937\pm 0.026$ $0.35^{+0.07}_{-0.05}$
Exponential Planck+DESI+PantheonPlus $0.21$ $-0.974\pm 0.028$ $0.27^{+0.10}_{-0.07}$
Planck+DESI+DESY5 $0.22$ $-0.942\pm 0.026$ $0.32^{+0.08}_{-0.05}$
JBP Planck+DESI+PantheonPlus $0.21$ $-0.985\pm 0.027$ $0.24^{+0.06}_{-0.04}$
$4.6$ $-0.988\pm 0.027$ $4.2\pm 0.9$
Planck+DESI+DESY5 $0.21$ $-0.945\pm 0.026$ $0.27^{+0.05}_{-0.03}$
$4.7$ $-0.946\pm 0.026$ $3.6\pm 0.5$
Logarithmic Planck+DESI+PantheonPlus $0.29$ $-0.979\pm 0.028$ $0.34^{+0.10}_{-0.07}$
Planck+DESI+DESY5 $0.26$ $-0.930^{+0.027}_{-0.026}$ $0.39^{+0.08}_{-0.05}$
BA Planck+DESI+PantheonPlus $0.28$ $-0.974^{+0.027}_{-0.028}$ $0.33^{+0.08}_{-0.06}$
Planck+DESI+DESY5 $0.28$ $-0.937^{+0.026}_{-0.027}$ $0.37^{+0.06}_{-0.04}$
Table 7: Constraints at 68% CL on the pivot redshift $z_{p}$, the corresponding value of the EoS $w(z_{p})$, and the redshift $z_{c}$ where the EoS crosses the phantom divide, for Planck+DESI+PantheonPlus and Planck+DESI+DESY5.
In this article, we have emphasized the resilience of the results recently delivered by the DESI collaboration, showing that DESI BAO data, in combination with Planck CMB observations and two different catalogs of SN distance moduli measurements (i.e., PantheonPlus and DESY5), consistently indicate a preference for DDE across various parameterizations of the EoS. While the primary goal was to confirm that this preference remains stable regardless of the specific DDE model, minor differences have emerged across the five cases analyzed, warranting further investigation. In this appendix, we explore these differences in more detail, aiming to provide a stronger physical interpretation of the results presented in the manuscript. For all five models, we reconstruct the evolution of the EoS with redshift, $w(z)$, based on the constraints on $w_{0}$ and $w_{a}$ inferred from the Planck+DESI+PantheonPlus and Planck+DESI+DESY5 datasets. In Fig. 7, we present the mean value of $w(z)$ (dashed lines), along with its uncertainties at the 68% (dark regions) and 95% (light regions) CL, across the redshift range $0\lesssim z\lesssim 6$ for all models and both data combinations. From the reconstructed shape of $w(z)$, we extract crucial information that helps compare the different models and clarify the results presented in the manuscript. Specifically, in Tab. 7, we present the results for:
  • (i)
    the pivot redshift $z_{p}$ and the corresponding value of the EoS, $w(z_{p})$, which indicate the redshift and the EoS value at which $w(z)$ is best constrained by the two datasets across the five models; [^13] See, e.g., Refs [397, 398] for discussions on the importance of the pivot redshift.
  • (ii)
    the redshift $z_{c}$ when the EoS crosses the phantom divide (i.e., $w(z_{c})=-1$), informing us of when the phantom crossing occurs, along with their respective uncertainties.
Starting from the baseline CPL model as the reference case, [^14] Note that the features presented in this appendix for the CPL model have been discussed in detail by the DESI collaboration – see, e.g., Sec. 5.2 of Ref. [336]. As we essentially recover all of the DESI results, we omit further discussion of the CPL model here. we summarize the main features and differences across the various models.
Figure 7: Evolution of $w(z)$ for $60<z<$ 6 across all DDE models, inferred from CMB+DESI+PantheonPlus (left panels) and CMB+DESI+DESY5 (right panels). The dashed lines represent the mean values, while the dark and light shaded regions indicate the $1\sigma$ and $2\sigma$ uncertainties, respectively.
  • Exponential: As seen when comparing the top panels with those in the second row from the top in Fig. 7, from a phenomenological perspective, the exponential parameterization closely resembles the CPL model. This similarity was already highlighted in the main text when comparing the improvement in $\Delta\chi^{2}$ over $\Lambda$CDM (which is quite similar for both scenarios). The agreement in predictions is further supported by Tab. 7. The only noticeable difference between the models is a slightly smaller pivot redshift $z_{p}$ in the exponential parameterization. However, the EoS at this pivot is constrained with comparable precision. Additionally, the predictions regarding the redshift of the phantom crossing $z_{c}$ agree within one standard deviation for both Planck+DESI+PantheonPlus and Planck+DESI+DESY5.
  • JBP: Among the five models analyzed, the JBP parameterization presents a more nuanced phenomenology regarding the evolution of the DE EoS. As shown in the third panels from the top in Fig. 7, due to its quadratic nature in the scale factor, the evolution of the EoS within the JBP parameterization crosses $w=-1$ twice. At very low redshift, it behaves similarly to the other parameterizations, remaining within the quintessence region $w(z)>-1$, albeit with larger uncertainties compared to the other models. The first quintessence-to-phantom transition is estimated to occur at $z=0.24^{+0.06}_{-0.04}$ ($z=0.27^{+0.05}_{-0.03}$) for Planck+DESI+PantheonPlus (Planck+DESI+DESY5) at 68% CL. After this transition, $w(z)$ approaches a minimum value around $z=1$ before rising towards less negative values. Eventually, a second phantom-to-quintessence crossing occurs at $z=4.2\pm 0.9$ ($z=3.6\pm 0.5$) for Planck+DESI+PantheonPlus (Planck+DESI+DESY5), both at 95% CL. This behavior contrasts with other models, where the EoS remains within the phantom regime, often trending towards more negative $w(z)$ values at high redshift. In contrast, within the JBP model, the EoS cannot move towards (more) phantom values but is compelled to transition back towards less negative values at $z>1$. The double crossing of regimes achieved within this parameterization is also reflected in the pivot scale $z_{p}$, at which the EoS is well constrained by data. In Tab. 7, we distinguish between two different regimes: the redshift range $10<z<$ 1 (capturing the first quintessence-to-phantom crossing) and the range $z>1$ (covering the second phantom-to-quintessence crossing). In these two regions, we identify two distinct pivot redshifts: the first at $z_{p}\sim 0.21$ and the second at $z_{p}\sim 4.6$. In both cases (and for both datasets), the EoS is constrained within the same minimal error. This confirms that, within this model, due to the functional form of the EoS, constraints at low redshift (i.e., around $z\sim 0.21$) also dictate the behavior of the parameterization at higher redshifts. The interplay between low and high redshift behaviors, as highlighted by the two pivot redshifts, could contribute to the increased uncertainties at low redshift and the tilting of the probability contours seen in Fig. 6. As discussed in the main manuscript, this model offers relatively modest improvements in the fit compared to $\Lambda$CDM, particularly in datasets covering $z\gtrsim 1$, where the model’s deviations from the others become more pronounced.
  • Logarithmic: When it comes to the logarithmic parameterization, the behavior of $w(z)$ for $z\lesssim 1$, shown in the fourth panel of Fig. 7, is similar to that of the CPL and exponential models. This is also reflected in the values we inferred for $z_{p}$, $w(z_{p})$, and $z_{c}$, all summarized in Tab. 7 and consistent with those models. However, we observe that at $z\gtrsim 1$, the EoS is forced down into deep phantom values, and the descent towards these very negative values is steeper than in the CPL and exponential cases. This is due to the fact that at $z\gtrsim 1$, the scale factor $a$ approaches small values (moving towards $a\to 0$), causing $\log(a)$ to decrease to negative values quite rapidly. This sudden decline in $w(z)$ for $z\gtrsim 1$ can lead to changes in the fit to data spanning $1\lesssim z\lesssim 3$, which is covered by BAO and SN observations, resulting in the differences in the $\chi^{2}$ of the fit discussed in the manuscript.
  • BA: Last but not least, the evolution of $w(z)$ obtained for the BA model is presented in the bottom panel of Fig. 7. As we argued in the manuscript, this model provides the most significant improvement in the fit over $\Lambda$CDM across all datasets analyzed in this study. Therefore, it is interesting to examine what is different in the evolution of the EoS compared to the other models. Looking at the low-redshift part of the EoS, we see that the model behaves very similarly to CPL (and its relatives). However, for the pivot redshift, we obtain $z_{p}=0.28$, slightly larger than in any other model, while $w(z_{p})$ takes values consistent with the other cases. We estimate the quintessence-to-phantom transition to occur at $z_{c}=0.33^{+0.08}_{-0.06}$ ($z_{c}=0.37^{+0.06}_{-0.04}$) for Planck+DESI+PantheonPlus (Planck+DESI+DESY5) at 68% CL. The most noticeable difference in the EoS occurs at $z\gtrsim 1$. In all other models studied so far, the EoS either moved deeply into phantom values (characterized by more or less steep functional forms of $w(z)$) or was compelled to increase back towards quintessence-like values in the JBP model. Referring to the bottom panel of Fig. 7, we observe that for $z\gtrsim 1$, the evolution of $w(z)$ in the BA model remains phantom but does not trend towards very negative values. Instead, $w(z)$ stabilizes on a sort of second plateau that is distinctive of the BA model.